how do I solve? 6x – 0.37 = 5(x – 0.28)
Answer by longjonsilver(2297)

6x – 0.37 = 5(x – 0.28)
6x – 0.37 = 5x – 1.4
x – 0.37 = – 1.4
x = – 1.03

jon